OpenStudy (anonymous):

The first four terms in a quadratic sequence are shown below. 6, 16, 32, 54, ... What is the next term in the sequence?

OpenStudy (king):

82

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

Thanks can you explain how you got that ?

OpenStudy (anonymous):

See, 16 - 6 = 10 32 - 16 = 16 54 - 32 = 22.. So, every number is increasing in terms of 6 like 10, 16, 22, so next increase will be of 28.. Therefore, 54 + 28 = 82 is the answer..

OpenStudy (king):

cuz in the first thing 6,16 to get the 2nd term the first term is added by 10 in the second one 16,32 to get 3rd term the 2nd term is added by 16 to get 4th term added by 22 so to get 5 th tern add by 28!

OpenStudy (anonymous):

ooh okay i get it @waterineyes i like the way you expained it thanks :)

OpenStudy (anonymous):

Welcome dear..